OpenStudy (dls):

Solve for b:

OpenStudy (dls):

\[\LARGE 9a^2 b^2=3a^2\]

OpenStudy (anonymous):

\[\huge 3\cancel{3a^2}b^2=\cancel{3a^2}\]

OpenStudy (dls):

how many values of b?

OpenStudy (skullpatrol):

How many values multiplied by themselves equal 1/3?

OpenStudy (dls):

?

OpenStudy (skullpatrol):

?

OpenStudy (dls):

what do u mean?

OpenStudy (skullpatrol):

I mean How many values multiplied by themselves equal 1/3?

OpenStudy (skullpatrol):

$$b^2=\frac{1}{3}$$

OpenStudy (anonymous):

you can also re-write as b^2-a^2=0 (b-a)(b+a)=0 difference of two squares hence two answeres

OpenStudy (anonymous):

\[b^2-\frac{1}{3}=0\] \[(b-\sqrt{1/3})(b+\sqrt{1/3})=0\]

OpenStudy (shubhamsrg):

You can not cancel out "a" since we don't know if that'll be equal to 0 or not. => a^2 ( (3b)^2 -1) = 0

OpenStudy (dls):

??

OpenStudy (dls):

\[\LARGE 9a^2 b^2-3a^2=0\] \[\LARGE 3a^2(3 b^2-1)=0\] a=0 b=0 and other two real values and a and b are those ellipse equation wale constant here..how can it be 0

OpenStudy (shubhamsrg):

So a is not 0, that means you can cancel it out

OpenStudy (shubhamsrg):

so b= +- sqrt(1/3)

OpenStudy (dls):

how do we know a is not 0

OpenStudy (shubhamsrg):

SInce it denotes length of major axis of ellipse!

OpenStudy (dls):

yeah so point ellipse :P ellipse doesn't exist

OpenStudy (dls):

its not wrong mathematically..

OpenStudy (shubhamsrg):

yep, mathematically its not, theoretically it is

OpenStudy (dls):

okay :o